OpenStudy (anonymous):

evaluate the following integral: 5x+1/(2x+1)(x-1) dx

myininaya (myininaya):

\[\int\limits_{}^{}(5x+\frac{1}{2x+1)(x-1)}) dx\]

myininaya (myininaya):

I assume you are having problems with the \[\int\limits_{}^{}\frac{1}{(2x+1)(x-1)} dx\] and not the \[\int\limits_{}^{}5x dx \]

myininaya (myininaya):

You should try partial fractions for the one with fractions :)

OpenStudy (anonymous):

okay thank you so much! :)